OpenStudy (anonymous):

The graph of a direct variation function passes through the origin. always sometimes never

jimthompson5910 (jim_thompson5910):

direct variation equations are of the form y = k*x k is some fixed number

jimthompson5910 (jim_thompson5910):

examples y = 2x, in this case, k = 2 y = (1/3)x, k is now k = 1/3

OpenStudy (anonymous):

So it would be sometimes then?

jimthompson5910 (jim_thompson5910):

one sec

jimthompson5910 (jim_thompson5910):

go here https://www.desmos.com/calculator and type in y = k*x into the box next to 1 to graph y = kx

jimthompson5910 (jim_thompson5910):

it will ask to make k a slider, do so then move k around to change the value of k

jimthompson5910 (jim_thompson5910):

you can even have it automatically animate for you

OpenStudy (anonymous):

Well, I don't really need to know how to find that, because i know how to do that, it is just the definition I am unsure of.

jimthompson5910 (jim_thompson5910):

you'll see it when you graph

OpenStudy (anonymous):

Ok. thank you.

jimthompson5910 (jim_thompson5910):

if you aren't sure on how to graph, you can go to my graph here https://www.desmos.com/calculator/djomyiw5n4 and everything is set up. Move the slider around for k or animate it to see what happens

jimthompson5910 (jim_thompson5910):

based on those graphs, what is the final answer?